A mass of 2.3 kg is initially moving at 5 m/s. A constant force of 15N is applied to the mass for 6 seconds. Find (a) the initial momentum, (b) final momentum.


1
Expert's answer
2021-12-22T14:11:35-0500

Given:

"m=2.3\:\rm kg"

"v_1=5\:\rm m/s"

"F=15\:\rm N"

"t=6\:\rm s"


(a) The initial momentum

"p_i=mv_1=2.3*5=11.5\:\rm kg*m/s"


(b) the final momentum

"p_f=p_i+Ft=11.5+15*6=101.5\:\rm kg*m/s"